OpenStudy (moongazer):

how do you solve for the unknown angles in these two equations? -20cos(alpha)-30cos(theta)+40=0 eq.1 +20sin(alpha)-30sin(theta) = 0 eq.2

OpenStudy (moongazer):

help :)

OpenStudy (asnaseer):

to start with you can divide both equations by 10 to simplify them

OpenStudy (asnaseer):

then use the eq 1 to get and expression for cos(alpha) and eq 2 to get and expression for sin(alpha)

OpenStudy (asnaseer):

finally use the identity:\[\cos^2(\alpha)+\sin^2(\alpha)=1\]to solve this

OpenStudy (michele_laino):

please, note that I can write your equations as below: \[2 \cos \alpha +3 \cos \theta =4,2 \sin \alpha-3 \sin \theta=0\] now, please square both sides of those equation, and then add them side by side:

OpenStudy (moongazer):

i'm trying to solve it :)

OpenStudy (michele_laino):

after my steps, you should get this condition: \[\cos (\alpha+\theta)=\frac{ 1 }{ 2 }\] please check!

OpenStudy (moongazer):

I got cos(alpha+theta) = 1/4

OpenStudy (michele_laino):

that's right! I have made an error

OpenStudy (asnaseer):

@moongazer - have you tried the method I suggested?

OpenStudy (asnaseer):

the question is asking you to find both alpha and theta

OpenStudy (moongazer):

@asnaseer not yet. I'll try it too

OpenStudy (michele_laino):

now we have a relation between alpha and theta, which we can use together other two equations

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{-20cos(\alpha)-30cos(\theta)+40=0 \\ 20sin(\alpha)-30sin(\theta) = 0\\~\\~\\~\\ 2cos(\alpha)+3cos(\theta)=4 \\~\\ 2sin(\alpha)-3sin(\theta) = 0\\~\\~\\~\\ 2cos(\alpha)=4-3cos(\theta)\\~\\ 2sin(\alpha) = 3sin(\theta)\\~\\~\\~\\ \text{squaring}\\~\\~\\~\\ 4cos^2(\alpha)=16-24cos(\theta)+9cos^2(\theta)---------\color{red}{(1)}\\~\\ 4sin^2(\alpha) =9sin^2(\theta)---------\color{red}{(2)}\\~\\}\end{align}\) add 1 and 2 this should give the answr quickly

OpenStudy (michele_laino):

multipling by sin theta your first equation, we get: \[2 \sin \alpha \sin \theta =3(\sin \theta)^{2}\] whereas multipling both sides of your second equation by cos theta, we find: \[2 \cos \alpha \cos \theta=4 \cos \theta -3(\cos \theta)\]

OpenStudy (moongazer):

I got the answer doing what asnaseer said: ) the answer is alpha=46.567 theta=28.955 thanks :)

OpenStudy (michele_laino):

oops...\[2 \cos \alpha \cos \theta =4 \cos \theta -3 (\cos \theta)^{2}\].

OpenStudy (moongazer):

@Michele_Laino I'll also try your solution. :) thanks

OpenStudy (asnaseer):

@moongazer - yw :)

OpenStudy (michele_laino):

substituting those above relation into the condition cos(Alpha+theta)=1/4, we get: \[3-4 \cos \theta =\frac{ 1 }{ 4 }\]

OpenStudy (michele_laino):

Thanks! @moongazer

OpenStudy (michele_laino):

oops...\[3-4 \cos \theta=\frac{ 1 }{ 2 }\]

OpenStudy (moongazer):

@Michele_Laino I think it should be: 1/2 = 4cos(theta) - 3 I also got the answer using your method. :) thanks!

OpenStudy (michele_laino):

yes! you are right, I'm very sorry!

OpenStudy (michele_laino):

thanks! again! @moongazer